# ClinKD: Cross-Modal Clinical Knowledge Distiller For Multi-Task Medical Images

### v3

@@ Abstract
Medical Visual Question Answering (Med-VQA) represents a critical and challenging subtask within the general VQA domain. Despite significant advancements in general Visual Question Answering (VQA), multimodal large language models (MLLMs) still exhibit substantial limitations when handling multi-task VQA scenarios. These limitations manifest through erroneous spatial localization and misinterpretation of medical images, which primarily arise from two fundamental issues: inadequate image-text alignment and insufficient medical knowledge in general-purpose MLLMs for specialized medical applications. To address these issues, we introduce the Cross-Modal Clinical Knowledge Distiller (ClinKD), an innovative framework designed to enhance image-text alignment and establish more effective medical knowledge adaptation mechanisms, which enables MLLMs to adapt to medical knowledge. Our extensive experimental evaluations demonstrate that the ClinKD achieves state-of-the-art performance on the Med-GRIT-270k dataset, a challenging medical benchmark containing fine-grained multi-task QA pairs. The results indicate that our approach not only significantly improves image-text alignment but also effectively enables MLLMs to adapt to the medical knowledge. The source code for ClinKD is available at: https://github.com/overloadedHenry/ClinKD. Refer to caption Figure 1: The left part shows that the BiRD [14] only utilizes traditional supervised fine-tuning (SFT) and 2D-RoPE [32] to obtain the capabilities of understanding and grounding medical images. The right part is our method which uses Med-CLIP Guided RoPE for enhancing image-text alignment and Pseudo-Labels Medical Knowledge Distillation for filling the gap of medical knowledge in ClinKD. During the SFT procedure, we will use Reflective Correction Training to enable answers with low semantic similarity or incorrect groundings to be reused.
@@ 1 Introduction
Medical Visual Question Answering (Med-VQA) [25, 1, 10] is a critical domain for applying MLLMs [37, 21, 4] to medical image analysis [19]. Recent advancements in MLLMs have enabled them to achieve preliminary capabilities in image analysis, annotation, and user instruction compliance [5]. Previous studies have made multiple attempts [38, 15, 11, 29, 18, 14] to enhance the performance of general-purpose MLLMs in Med-VQA.
@@
For instance, LLaVA-Med [18] extends instruction-tuning techniques for MLLMs to the medical domain. This work applied instruction tuning and curated image-text datasets to align visual-textual features in LLaVA [27], enabling LLaVA-Med to recognize and analyze medical images. However, this system lacked fine-grained image understanding, causing erroneous spatial localization when questioned about specific pathological locations [9, 14].
@@
Subsequently, the BiRD [14], a model that incorporated fine-grained image recognition by developing a biomedical refer-and-ground instruction-tuning dataset, addresses the absence of grounding capabilities [40, 42] in LLaVA-Med’s work [18]. Nevertheless, as shown in Figure 1, due to its reliance on traditional supervised fine-tuning (instruction-tuning) [28, 17, 45] and conventional 2D Rotary Position Embedding (RoPE) [32], BiRD tends to generate erroneous spatial localization and misinterpretation of medical images because of incomplete image-text alignment [35, 43]. Moreover, the visual encoder of BiRD is constrained by the lack of medical knowledge, resulting in difficulties in medical knowledge adaptation [47, 24].
@@
To overcome these limitations, we propose a medical distillation framework named ClinKD to boost capabilities in the Med-VQA domain. Specifically, our method mainly focuses on two issues: incomplete image-text alignment and the professional medical knowledge gap between the general-purpose MLLMs and medical domain that requires finer-grained medical knowledge [22, 41].
@@
For the first issue, we propose Med-CLIP Guided Rotary Position Embedding (MCG-RoPE), a novel position embedding method that focuses on the inter-modal and intra-modal intervals during joint image-text training, while ensuring the equivalence of context intervals immediately adjacent to images. Compared to traditional RoPE [32], the MCG-RoPE uses the distinct index intervals to better capture cross-modal information, thereby achieving more effective image-text alignment.
@@
For the second issue, we propose Clinical Knowledge Distiller (ClinKD) which consists of Pseudo-Labels Medical Knowledge Distillation and Reflective Correction Training. Pseudo-Labels Medical Knowledge Distillation provides prior medical knowledge by using pseudo-labels [20, 6], filling the gap of prior medical knowledge so that the model can better adapt to medical knowledge during the supervised fine-tuning. Reflective Correction Training allows samples with low semantic similarity to be sent for training again after being enhanced by MLLMs.
@@
  • We introduce the ClinKD, an innovative distiller designed to establish more effective medical knowledge adaptation mechanisms.
@@
  • For bridging medical knowledge gap between general-purpose MLLMs and specialized medical applications, we propose the Pseudo-Labels Medical Knowledge Distillation, filling the medical knowledge gap before supervised fine-tuning.
  • Extensive experimental results show that the proposed ClinKD achieves the best performance on Med-GRIT-270k [14].
@@ 2.1 Biomedical Instruction-Tuning Datasets
Many Med-VQA datasets such as VQA-RAD [16] and SLAKE [26] contain only medical images and simple QA pairs. Specifically, these QA pairs do not have the fine-grained descriptions of pathological features. This limitation constrains the ability of MLLMs to perform accurate spatial localization. To boost and evaluate MLLMs’ capabilities in the Med-VQA domain, many biomedical instruction-tuning datasets are designed.
For example, the LLaVA-Med-qa0.2k [46] contains 50 medical images sampled from several modalities such as CT, Angiography and PET and 193 questions according to these medical images. The medical QA pairs are generated by GPT-4 based on the metadata. This dataset is designed to evaluate the capabilities of describing medical images, specifically the biomedical features.
Unlike the LLaVA-Med-qa0.2k [46], the Med-GRIT-270k [14] includes multi-task fine-grained medical QA pairs especially for grounding spatial locations. It is designed to enhance the model’s capabilities in referring and grounding. This multi-task dataset comprises four distinct tasks: Visual Grounding (VG), Referring Object Classification (ROC), Referring Captioning (RC), and Medical Image Analysis (MIA). The VG task evaluates the model’s capability of accurately matching text descriptions to corresponding image regions. The RC task assesses the model’s capability to recognize specific image areas and generate descriptive captions for them. The ROC task examines the model’s understanding of textual information related to particular image regions and their associated visual details. The MIA task evaluates the model’s comprehension of medical images and their multi-modal context. All data is sourced from eight modalities: CT, MR, X-ray, PET, Endoscopy, Dermoscopy, Fundus, and Ultrasound. Utilizing data from these eight modalities allows for a more comprehensive evaluation of the model’s capabilities.
@@ 2.2 Multimodal Large Language Models (MLLMs) for Biomedicine
With the development of MLLMs, an increasing number of approaches have been applied to the medical field, such as BioMedGPT [44], LLaVA-Med [18], and BiRD [14]. These methods drive the development of MLLMs in the biomedical domain from different perspectives. For instance, LLaVA-Med [18] applied instruction fine-tuning to the medical imaging field and built a medical question-answering system based on LLaVA [27], which provided reasonable responses by processing user-submitted images and instructions. However, LLaVA-Med did not support fine-grained interactions with medical images, meaning that the system cannot precisely locate pathological regions within the images. To address the lack of referring and grounding capabilities, the BiRD [14], which is based on Qwen-VL [3], leveraged a finer-grained medical dataset. However, the ViT of BiRD was not trained by sufficient medical data before being frozen, resulting in a gap in medical knowledge domain. Compared with their work [14, 18], our method modifies the position embedding method and introduces an effective pseudo-labels medical knowledge distillation framework.
@@ 2.3 Rotary Position Embedding (RoPE)
The introduced RoPE [32] makes the relative position features of the context in the text more easily captured by LLMs. Further, the RoPE-Mixed [12] introduced mixed learnable frequencies to RoPE, enabling it to handle diagonal directions and making RoPE itself learnable. It showed more efficiency when dealing with diagonal features of the image. The VideoRoPE [36] extended the diagonal layout and variable frequency to the video domain, demonstrating its advantages in video analysis tasks. Unlike existing methods, our approach focuses on 2D multi-task medical images where capturing semantic information is challenging. We utilize cross-modal intervals to enable the model to better distinguish modal information, thereby improving image-text alignment.
@@ 3 Methodology
As shown in Figure 2, we aim at enhancing the ClinKD’s performance by mainly leveraging Med-CLIP Guided Rotary Position Embedding, Pseudo-Labels Medical Knowledge Distillation, Reflective Correction Training and Semantic-Aware Selective Generation. The proposed Med-CLIP Guided Rotary Position uses the distinct inter-modal and intra-modal space to improve image-text alignment. The distiller provides prior medical knowledge to enhance the model’s capability of medical knowledge adaptation. In the inference procedure, the Semantic-Aware Selective Generation will allow the model to choose the answer with the best score.
@@ 3.1 Med-CLIP Guided Rotary Position Embedding
The 2D Rotary Position Embedding (RoPE) [32] tends to lack attention on cross-modal features, so we propose the Med-CLIP Guided Rotary Position Embedding (MCG-RoPE) to modify rotation angles that are crucial for image-text alignment. Figure 3 shows the example of our method.
Distinct Inter-Modal Intervals. Suppose we have a sentence-image-sentence position index sequence:
$S_{idx}=($ $T_{1},T_{2}\cdots,T_{\ell},V_{\ell+1},V_{\ell+2},\cdots,$ (1)
$V_{\ell+wh},T_{\ell+wh+1},T_{\ell+wh+2},\cdots,T_{\ell+wh+k})$
where $T_{i}=(t_{i},t_{i})$ denotes the 2D position index of text tokens, $V_{i}=(v^{(1)}_{i},v^{(2)}_{i})$ denotes the 2D image index.
To distinguish between intra-modal features and inter-modal features, the indexes of image tokens need to be transformed.
$\left(v_{i}^{(1)},v_{i}^{(2)}\right)\xrightarrow{\mathcal{F}}\left(\alpha_{1}v_{i}^{(1)}+\lambda_{1},\alpha_{2}v_{i}^{(2)}+\lambda_{2}\right)=\mathcal{V}_{i}$ (2)
where $\mathcal{V}_{i}$ is the position after being scaled and shifted. The $\lambda_{1},\lambda_{2},\alpha_{1},\alpha_{2}$ are parameters. The $\mathcal{F}$ denotes the operator of shifting and scaling.
The inter-modal intervals should be consistent, and thus the following equation needs to be satisfied:
$\begin{cases}\alpha_{1}+\lambda_{1}-\ell=\ell+wh+1-(\alpha_{1}h+\lambda_{1})\\ \alpha_{2}+\lambda_{2}-\ell=\ell+wh+1-(\alpha_{2}w+\lambda_{2})\end{cases}$ (3)
We need to fix $\alpha_{1},\alpha_{2}$ first to set the intervals between each image’s position index. So the general solution is
$\begin{cases}\lambda_{1}=\ell+\frac{[wh+1-\alpha_{1}(h+1)]}{2}\\ \lambda_{2}=\ell+\frac{[wh+1-\alpha_{2}(w+1)]}{2}\end{cases}$ (4)
After setting intervals, the new sequence can be expressed as:
$S_{h-idx}=($ $t_{1},\cdots,t_{\ell},\alpha_{1}v^{(1)}_{\ell+1}+\lambda_{1},\cdots,$ (5)
$\alpha_{1}v^{(1)}_{\ell+wh}+\lambda_{1},,t_{\ell+wh+1},\cdots,t_{\ell+wh+k})$
where $S_{h-idx}$ is the position index on height dimension,
$S_{w-idx}=($ $t_{1},\cdots,t_{\ell},\alpha_{2}v^{(2)}_{\ell+1}+\lambda_{2},\cdots,$ (6)
$\alpha_{2}v^{(2)}_{\ell+wh}+\lambda_{2},,t_{\ell+wh+1},\cdots,t_{\ell+wh+k})$
where $S_{w-idx}$ is the position index on width dimension.
Distinct Interval-Based Rotation. Suppose we have a query/key vector: $q\in\mathbb{R}^{d_{h}+d_{w}}$ where $\mathbb{R}^{d_{h}+d_{w}}$ denotes the $d_{h}+d_{w}$ dimension Euclidean space. The $d_{h}$, $d_{w}$ denotes the embedding dimension on height dimension and width dimension respectively. Then $q$ will be split into $q^{(h)}\in\mathbb{R}^{d_{h}}$ and $q^{(w)}\in\mathbb{R}^{d_{w}}$
The corresponding frequency on the two dimensions is set as:
$\omega_{i}^{(h)}=10000^{-\frac{2i}{d_{h}}}$ (7)
$\omega_{i}^{(w)}=10000^{-\frac{2i}{d_{w}}}$
Then the angles of rotation are expressed as:
$\theta_{x_{h},i}^{(h)}=x_{h}\cdot\omega_{i}^{(h)}$ (8)
$\theta_{x_{w},i}^{(w)}=x_{w}\cdot\omega_{i}^{(w)}$
where $x_{h}$, $x_{w}$ denote the transformed position index on height and width dimension respectively. This means the angle of rotation will be determined by inter-modal and intra-modal intervals. A large rotation angle brings information of modal transformation while a smaller angle contains fine-grained features of intra-modal. The rotation matrix can be defined as:
$\boldsymbol{R}_{\boldsymbol{\Theta},x_{h}}^{d_{h}}=\begin{pmatrix}R_{\theta_{x_{h},1}^{(h)}}&&\\ &\ddots&\\ &&R_{\theta_{x_{h},d_{h}/2}^{(h)}}\end{pmatrix}$ (9)
$\boldsymbol{R}_{\boldsymbol{\Theta},x_{w}}^{d_{w}}=\begin{pmatrix}R_{\theta_{x_{w},1}^{(w)}}&&\\ &\ddots&\\ &&R_{\theta_{x_{w},d_{w}/2}^{(w)}}\end{pmatrix}$
where $R_{\theta_{x_{h},i}^{(h)}}$ and $R_{\theta_{x_{w},i}^{(w)}}$ are defined as:
$R_{\theta_{x_{w},i}^{(w)}}=\begin{pmatrix}{\operatorname{cos}\theta_{x_{w},i}^{(w)}}&{-\operatorname{sin}\theta_{x_{w},i}^{(w)}}\\ {\operatorname{sin}\theta_{x_{w},i}^{(w)}}&{\operatorname{cos}\theta_{x_{w},i}^{(w)}}\\ \end{pmatrix}$ (10)
$R_{\theta_{x_{h},i}^{(h)}}=\begin{pmatrix}{\operatorname{cos}\theta_{x_{h},i}^{(h)}}&{-\operatorname{sin}\theta_{x_{h},i}^{(h)}}\\ {\operatorname{sin}\theta_{x_{h},i}^{(h)}}&{\operatorname{cos}\theta_{x_{h},i}^{(h)}}\\ \end{pmatrix}$
The rotated vector is shown in Eq. (11) :
$\widetilde{q}=\left[\boldsymbol{R}_{\boldsymbol{\Theta},x_{h}}^{d_{h}}{q^{(h)}};\boldsymbol{R}_{\boldsymbol{\Theta},x_{h}}^{d_{h}}{q^{(w)}}\right]$ (11)
3.2 Pseudo-Labels Medical Knowledge Distillation (Pseudo-KD)
Our clinical knowledge distiller enables the student model to learn the output distribution of the teacher model. Therefore, we provide prior knowledge of medical images to the student model through Pseudo-Labels Medical Knowledge Distillation.
The distillation loss function is defined by a linear combination of cross-entropy loss and the Kullback-Leibler Divergence (KL Divergence).
The cross-entropy loss is defined as:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}y_{i}\log p_{s}(z_{s_{i}})$ (12)
where $\mathcal{L}_{CE}$ denotes the cross-entropy loss, $p_{s}(z_{i})$ is the logits generated by the student ClinKD model and $y_{i}$ is the hard label of ground truth.
The KL Divergence mentioned before is defined as:
$\mathcal{L}_{KD}=T^{2}\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})$ (13)
where $\mathcal{L}_{KD}$ denotes the loss function constructed by KL Divergence, $\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})$ is the KL Divergence, $T$ is called temperature which can smooth the distribution and the $\epsilon_{s_{j}}$, $\epsilon_{t_{j}}$ are noise values.
Furthermore, the $\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})$ is:
$\mathbb{D}^{\boldsymbol{\epsilon_{t}},\boldsymbol{\epsilon_{s}}}_{KL}(p_{t}\|p_{s})=\sum_{i=1}^{n}p_{t}(z_{t_{i}}+\epsilon_{t_{j}})\log\frac{p_{t}(z_{t_{i}}+\epsilon_{t_{j}})}{p_{s}(z_{s_{i}}+\epsilon_{s_{j}})}$ (14)
where $p_{t}(z_{i})$ is the probability distribution generated by the teacher ClinKD model, $p_{s}(z_{i})$ denotes the probability distribution generated by the student ClinKD model.
For details, the distributions are smoothed by temperature $T$:
$p_{s}(z_{s_{j}}+\epsilon_{s_{j}})=\dfrac{exp((z_{s_{j}}+\epsilon_{s_{j}})/T)}{\displaystyle\sum_{i}^{n}exp((z_{s_{i}}+\epsilon_{s_{i}})/T)}$ (15)
$p_{t}(z_{t_{j}}+\epsilon_{t_{j}})=\dfrac{exp((z_{t_{j}}+\epsilon_{t_{j}})/T)}{\displaystyle\sum_{i}^{n}exp((z_{t_{i}}+\epsilon_{t_{i}})/T)}$
Finally the whole distillation loss is expressed as:
$\mathcal{L}_{Distill}=\dfrac{\mathcal{L}_{CE}+\alpha\mathcal{L}_{KD}}{1+\alpha}$ (16)
After obtaining teacher’s prior knowledge about medical images, the student ClinKD model will be trained by supervised fine-tuning (SFT).

### v4

@@ Abstract
Medical Visual Question Answering (Med-VQA) represents a critical and challenging subtask within the general VQA domain. Despite significant advancements in general VQA, multimodal large language models (MLLMs) still exhibit substantial limitations when handling multi-task VQA scenarios. These limitations manifest through erroneous spatial localization and misinterpretation of medical images, which primarily arise from two fundamental issues: inadequate image-text alignment and insufficient domain-specified knowledge for medical applications. To address these issues, we introduce the Cross-Modal Clinical Knowledge Distiller (ClinKD), an innovative framework designed to enhance image-text alignment and establish more effective medical knowledge transformation mechanisms, which enables MLLMs to perform better even when lacking prior medical knowledge. Our extensive experimental evaluations demonstrate that the ClinKD achieves state-of-the-art performance on several datasets which are challenging for the Med-VQA task. The results indicate that our approach not only significantly improves image-text alignment but also effectively enables MLLMs to adapt to the medical knowledge. The source code for ClinKD is available at: https://github.com/overloadedHenry/ClinKD. Refer to caption Figure 1: The left part shows that the BiRD huang2024BiRD only utilizes traditional supervised fine-tuning (SFT) and 2D-RoPE su2023roformerenhancedtransformerrotary to obtain the capabilities of understanding and grounding medical images. The right part is our method which uses Med-CLIP Guided RoPE for enhancing image-text alignment and Adaptive Confidence-Margin Curriculum Pseudo-KD for filling the gap of medical knowledge in ClinKD. During the SFT procedure, we will use Reflective Correction Training to enable answers with low semantic similarity or incorrect groundings to be reused.
@@ 1 Introduction
Medical Visual Question Answering (Med-VQA) Lin_2023 ; Abacha2019VQAMedOO ; Hasan2018OverviewOI is a critical domain for applying MLLMs yang2024modelmergingllmsmllms ; liang2024comprehensivesurveyguidemultimodal ; caffagni-etal-2024-revolution to medical image analysis li2024towards . Recent advancements in MLLMs have enabled them to achieve preliminary capabilities in image analysis, annotation, and user instruction compliance chen2025rllavaimprovingmedvqaunderstanding . Previous studies have made multiple attempts yang2025llmmedqaenhancingmedicalquestion ; kumar2024medvisionllamaleveragingpretrainedlarge ; he2024parameterefficientfinetuningmedicalmultimodal ; liu2024hcllmhistoricalconstrainedlargelanguage ; li2024llava ; huang2024BiRD to enhance the performance of general-purpose MLLMs in Med-VQA.
@@
For instance, LLaVA-Med li2024llava extends instruction-tuning techniques for MLLMs to the medical domain. This work applied instruction tuning and curated image-text datasets to align visual-textual features in LLaVA llava , enabling LLaVA-Med to recognize and analyze medical images. However, this system lacked fine-grained image understanding, causing erroneous spatial localization when questioned about specific pathological locations gai2024enhancing ; huang2024BiRD .
@@
Subsequently, the BiRD huang2024BiRD , a model that incorporated fine-grained image recognition by developing a biomedical refer-and-ground instruction-tuning dataset, addresses the absence of grounding capabilities you2023ferretrefergroundgranularity ; zhang2024ferretv2improvedbaselinereferring in LLaVA-Med’s work li2024llava . Nevertheless, as shown in Figure 1, due to its reliance on traditional supervised fine-tuning (instruction-tuning) sft1 ; sft2 ; sft3 and conventional 2D Rotary Position Embedding (RoPE) su2023roformerenhancedtransformerrotary , BiRD tends to generate erroneous spatial localization and misinterpretation of medical images because of incomplete image-text alignment alignment1 ; alignment2 . Moreover, the visual encoder of BiRD is constrained by the lack of medical knowledge, resulting in difficulties in medical knowledge adaptation adaptation ; lin2025healthgptmedicallargevisionlanguage .
@@
To overcome these limitations, we propose a medical distillation framework named ClinKD to boost capabilities in the Med-VQA domain. Specifically, our method mainly focuses on two issues: incomplete image-text alignment and the professional medical knowledge gap between the general-purpose MLLMs and medical domain that requires finer-grained medical knowledge medical_fine-grained ; zeng2024visualorientedfinegrainedknowledgeediting .
@@
For the first issue, we propose Med-CLIP Guided Rotary Position Embedding (MCG-RoPE), a novel position embedding method that focuses on the inter-modal and intra-modal intervals during joint image-text training, while ensuring the equivalence of context intervals immediately adjacent to images. Compared to traditional RoPE su2023roformerenhancedtransformerrotary , the MCG-RoPE uses the distinct index intervals to better capture cross-modal information, thereby achieving more effective image-text alignment.
@@
For the second issue, we propose Clinical Knowledge Distiller (ClinKD) which consists of Adaptive Confidence-Margin Curriculum Pseudo-KD (Pseudo-KD) and Reflective Correction Training. Pseudo-KD provides prior medical knowledge by using pseudo-labels pseudolabels ; chen2023mixedpseudolabelssemisupervised , filling the gap of prior medical knowledge so that the model can better adapt to medical knowledge during the supervised fine-tuning. Reflective Correction Training allows samples with low semantic similarity to be sent for training again after being enhanced by MLLMs.
@@
  • We introduce the ClinKD, an innovative distiller designed to establish more effective medical knowledge adaptation mechanisms. The code and checkpoint will be released.
@@
  • For bridging medical knowledge gap between general-purpose MLLMs and specialized medical applications, we propose the Adaptive Confidence-Margin Curriculum Pseudo-KD, filling the medical knowledge gap before supervised fine-tuning.
  • Extensive experimental results show that the proposed ClinKD achieves competitive or superior performance on test set of diverse tasks and datasets.
@@ 2.1 Biomedical Instruction-Tuning Datasets
Many Med-VQA datasets such as VQA-RAD lau2018dataset and SLAKE liu2021slake contain only medical images and simple QA pairs. Specifically, these QA pairs do not have the fine-grained descriptions of pathological features. This limitation constrains the ability of MLLMs to perform accurate spatial localization. To boost and evaluate MLLMs’ capabilities in the Med-VQA domain, many biomedical instruction-tuning datasets are designed.
For example, the LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation contains 50 medical images sampled from several modalities such as CT, Angiography and PET and 193 questions according to these medical images. The medical QA pairs are generated by GPT-4 based on the metadata. This dataset is designed to evaluate the capabilities of describing medical images, specifically the biomedical features.
Unlike the LLaVA-Med-qa0.2k zhang2025biomedclipmultimodalbiomedicalfoundation , the Med-GRIT-270k huang2024BiRD includes multi-task fine-grained medical QA pairs especially for grounding spatial locations. It is designed to enhance the model’s capabilities in referring and grounding. This multi-task dataset comprises four distinct tasks: Visual Grounding (VG), Referring Object Classification (ROC), Referring Captioning (RC), and Medical Image Analysis (MIA). The VG task evaluates the model’s capability of accurately matching text descriptions to corresponding image regions. The RC task assesses the model’s capability to recognize specific image areas and generate descriptive captions for them. The ROC task examines the model’s understanding of textual information related to particular image regions and their associated visual details. The MIA task evaluates the model’s comprehension of medical images and their multi-modal context. All data is sourced from eight modalities: CT, MR, X-ray, PET, Endoscopy, Dermoscopy, Fundus, and Ultrasound. Utilizing data from these eight modalities allows for a more comprehensive evaluation of the model’s capabilities.
@@ 2.2 Multimodal Large Language Models (MLLMs) for Biomedicine
With the development of MLLMs, an increasing number of approaches have been applied to the medical field, such as BioMedGPT Zhang_2024 , LLaVA-Med li2024llava , and BiRD huang2024BiRD . These methods drive the development of MLLMs in the biomedical domain from different perspectives. For instance, LLaVA-Med li2024llava applied instruction fine-tuning to the medical imaging field and built a medical question-answering system based on LLaVA llava , which provided reasonable responses by processing user-submitted images and instructions. However, LLaVA-Med did not support fine-grained interactions with medical images, meaning that the system cannot precisely locate pathological regions within the images. To address the lack of referring and grounding capabilities, the BiRD huang2024BiRD , which is based on Qwen-VL bai2023qwenvlversatilevisionlanguagemodel , leveraged a finer-grained medical dataset. However, the ViT of BiRD was not trained by sufficient medical data before being frozen, resulting in a gap in medical knowledge domain. Compared with their work huang2024BiRD ; li2024llava , our method modifies the position embedding method and introduces an effective Confidence-Margin Curriculum Pseudo-KD framework.
@@ 2.3 Rotary Position Embedding (RoPE)
The introduced RoPE su2023roformerenhancedtransformerrotary makes the relative position features of the context in the text more easily captured by LLMs. Further, the RoPE-Mixed heo2024rotarypositionembeddingvision introduced mixed learnable frequencies to RoPE, enabling it to handle diagonal directions and making RoPE itself learnable. It showed more efficiency when dealing with diagonal features of the image. The VideoRoPE wei2025videoropemakesgoodvideo extended the diagonal layout and variable frequency to the video domain, demonstrating its advantages in video analysis tasks. Unlike existing methods, our approach focuses on 2D multi-task medical images where capturing semantic information is challenging. We utilize cross-modal intervals to enable the model to better distinguish modal information, thereby improving image-text alignment.
@@ 3 Methodology
As shown in Figure 2, we aim at enhancing the ClinKD’s performance by mainly leveraging Med-CLIP Guided Rotary Position Embedding, Pseudo-KD, Reflective Correction Training and Semantic-Aware Selective Generation. The proposed Med-CLIP Guided Rotary Position uses the distinct inter-modal and intra-modal space to improve image-text alignment. The distiller provides prior medical knowledge to enhance the model’s capability of medical knowledge adaptation. In the inference procedure, the Semantic-Aware Selective Generation will allow the model to choose the answer with the best score.
@@ 3.1 Med-CLIP Guided Rotary Position Embedding
The standard 2D Rotary Position Embedding (RoPE) su2023roformerenhancedtransformerrotary does not explicitly account for cross-modal relationships between image and text tokens. To address this, we introduce the Med-CLIP Guided Rotary Position Embedding (MCG-RoPE), which deal with tokens from different modals by considering splited queries to improve image-text alignment. Figure 3 illustrates an example of our approach.
Distinct Inter-Modal Intervals.
Consider a sequence of position indices for a text–image–text pattern:
$S_{\mathrm{idx}}=\bigl{(}T_{1},T_{2},\dots,T_{\ell},\;V_{\ell+1},\dots,V_{\ell+wh},\;T_{\ell+wh+1},\dots,T_{\ell+wh+k}\bigr{)},$ (1)
where each text index $T_{i}=(t_{i},t_{i})\in\mathbb{R}^{2}$ and each image index $V_{j}=(v^{(1)}_{j},v^{(2)}_{j})\in\mathbb{R}^{2}$. We apply an affine transform $\mathcal{F}$ to each image index so that intra-modal and inter-modal gaps become equal along both axes:
$\mathcal{F}\bigl{(}v_{i}^{(1)},v_{i}^{(2)}\bigr{)}=\bigl{(}\alpha_{1}\,v_{i}^{(1)}+\lambda_{1},\;\alpha_{2}\,v_{i}^{(2)}+\lambda_{2}\bigr{)}=:\mathcal{V}_{i},$ (2)
with scaling factors $\alpha_{1},\alpha_{2}>0$ and offsets $\lambda_{1},\lambda_{2}\in\mathbb{R}$. By enforcing a pronounced interval difference between text and image tokens, the model is guided to focus on fine-grained visual details in medical images while maintaining clear separation from surrounding text.
Distinct Interval-Based Rotation.
Let $q\in\mathbb{R}^{d_{h}+d_{w}}$ be a query/key vector, split as
$q=\bigl{[}q^{(h)};q^{(w)}\bigr{]},\quad q^{(h)}\in\mathbb{R}^{d_{h}},\;q^{(w)}\in\mathbb{R}^{d_{w}}.$ (3)
By processing the height and width dimensions separately, the model can capture both coarse inter-modal alignment and fine intra-modal structure.
Define frequency scales for each half of the embedding:
$\omega_{i}^{(h)}=10\,000^{-\frac{2i}{d_{h}}},\quad\omega_{i}^{(w)}=10\,000^{-\frac{2i}{d_{w}}},\quad i=1,\dots,\tfrac{d_{h}}{2}\;\text{or}\;\tfrac{d_{w}}{2}.$ (4)
Given the transformed positions $x_{h}$ and $x_{w}$, the rotation angles satisfy
$\theta_{x_{h},i}^{(h)}=x_{h}\,\omega_{i}^{(h)},\quad\theta_{x_{w},i}^{(w)}=x_{w}\,\omega_{i}^{(w)}.$ (5)
We construct block-diagonal rotation matrices
$\mathbf{R}_{\Theta,x_{h}}^{d_{h}}=\mathrm{diag}\bigl{(}R_{\theta_{x_{h},1}^{(h)}},\dots,R_{\theta_{x_{h},\frac{d_{h}}{2}}^{(h)}}\bigr{)},\quad\mathbf{R}_{\Theta,x_{w}}^{d_{w}}=\mathrm{diag}\bigl{(}R_{\theta_{x_{w},1}^{(w)}},\dots,R_{\theta_{x_{w},\frac{d_{w}}{2}}^{(w)}}\bigr{)},$ (6)
where each $2\times 2$ block is
$R_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}.$ (7)
Finally, the transformed vector is
$\widetilde{q}=\bigl{[}\mathbf{R}_{\Theta,x_{h}}^{d_{h}}\,q^{(h)};\,\mathbf{R}_{\Theta,x_{w}}^{d_{w}}\,q^{(w)}\bigr{]}.$ (8)
By emphasizing distinct interval gaps and splitting the query into separate height and width components, MCG-RoPE enables the model to more effectively capture fine-grained features in medical images and achieve precise image–text alignment. The difference between 2D-RoPE and MCG-RoPE is shown in Appendix.
3.2 Adaptive Confidence-Margin Curriculum Pseudo-KD
In order to mitigate the negative impact of noisy pseudo-labels and more effectively guide the student model, we propose an Adaptive Confidence–Margin Curriculum Pseudo-KD strategy. This approach extends standard pseudo-label distillation by (1) weighting each sample’s distillation loss according to both the teacher’s confidence and the margin between top logits, and (2) employing a simple curriculum learning schedule that gradually relaxes the confidence threshold over training.
Confidence and Margin Computation.
Given an input example $x$, the teacher network produces logits $z_{t}=f_{t}(x)$. We compute the softmax probabilities
$p_{t_{i}}(x)\;=\;\frac{\exp\bigl{(}z_{t_{i}}\bigr{)}}{\sum_{j}\exp\bigl{(}z_{t_{j}}\bigr{)}}\,.$ (9)
We define the confidence and margin measures as
$C(x)\;=\;\max_{i}p_{t_{i}}(x)\quad,\quad\Delta(x)\;=\;p_{t_{(1)}}(x)\;-\;p_{t_{(2)}}(x)\,,$ (10)
where $p_{(1)}$ and $p_{(2)}$ denote the largest and second-largest probabilities, respectively.
Curriculum Threshold Schedule.
Let $t\in[0,1]$ denote the normalized training progress (i.e., fraction of total steps completed). We set an initial confidence threshold $\tau_{0}$ and a minimum threshold $\tau_{\min}$. The threshold decays linearly:
$\tau(t)\;=\;\tau_{\min}\;+\;(\tau_{0}-\tau_{\min})\,\bigl{(}1-t\bigr{)}\,.$ (11)
Adaptive Weighting Function.
We combine confidence and margin into a per-sample weight:
$w(x)\;=\;\begin{cases}\bigl{[}C(x)\bigr{]}^{\gamma}\,\bigl{[}\Delta(x)\bigr{]}^{\beta},&C(x)\;\geq\;\tau(t)\,,\\ 0,&\text{otherwise},\end{cases}$ (12)
where $\gamma,\beta\in(0,1]$ control sensitivity to confidence and margin.
Weighted Distillation Loss.
With student logits $z_{s}$ and temperature $T$, the weighted KL-divergence distillation loss is
$\mathcal{L}_{\mathrm{KD}}=\frac{1}{N}\sum_{i=1}^{N}w(x_{i})\;\mathbb{D}_{KL}\Bigl{(}softmax\bigl{(}z_{t_{i}}/T\bigr{)}\,\big{\|}\,softmax\bigl{(}z_{s_{i}}/T\bigr{)}\Bigr{)}\,.$ (13)
Overall Objective.
We combine the standard cross-entropy loss $L_{\mathrm{CE}}$ with our adaptive pseudo-KD term:
$\mathcal{L}_{Distill}\;=\;\dfrac{\mathcal{L}_{\mathrm{CE}}\;+\;\alpha\,\mathcal{L}_{\mathrm{KD}}}{1+\alpha}\quad,\quad\alpha\in[0,1]\,.$ (14)